# Real-World Remote Sensing Image Dehazing: Benchmark and Baseline

### v1

@@ IV-B Haze Density Estimation
where $I$ is the hazy image, $A$ is the atmospheric light, and $\omega$ is a constant (typically 0.95) to retain subtle haze effects. For each image, we calculated the mean dark channel value as a proxy for haze density. Statistical analysis across the dataset established thresholds at 110.58 for thin haze and 159.31 for thick haze, as depicted in Fig. 6. Images were accordingly categorized into thin (below 110.58), moderate (110.58–159.31), and thick (above 159.31) haze levels.
Figure 6: Haze density distribution across the RRSHID dataset.
@@ IV-C Dataset Characteristics
RRSHID encompasses 3,053 pairs of hazy and haze-free RS images, constituting the largest real-world dataset of its kind. These pairs are stratified by haze density into three categories, as detailed in Table I.
· Natural variations in illumination and their influence on haze perception.
These properties render RRSHID a crucial resource for developing robust RSID algorithms, assessing haze impacts on applications such as object detection [50, 51] and change detection [52, 53], and deepening insights into atmospheric effects on satellite imagery.
TABLE I: Distribution of Image Pairs in the RRSHID Dataset
@@ V-A1 Datasets
MCAF-Net is evaluated on the RRSHID dataset and four publicly available benchmarks: RSID [1], RICE1 [11], RICE2 [11] and StateHaze1K-thick [49]. These datasets are selected for their diverse characteristics:
@@
The RRSHID dataset, designed to reflect real-world RS challenges, complements these synthetic benchmarks to assess generalization.
@@ V-A3 Evaluation Metrics
To assess the performance on supervised hazy datasets, we employed three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Mean Squared Error (MSE). Higher PSNR and SSIM values signify improved image quality, while lower MSE indicates enhanced dehazing efficacy.
@@ V-B Experimental Results
TABLE II: Performance comparison of different methods on RRSHID dataset across various haze levels. Values in red and blue denote the best and second-best performance, respectively.
@@
PSNR SSIM MSE PSNR SSIM MSE PSNR SSIM MSE PSNR SSIM MSE
DCP[8] 18.46 0.4564 0.0192 17.80 0.4856 0.0238 18.39 0.4843 0.0208 18.22 0.4754 0.0213
FFA-Net [12] 17.08 0.4452 0.0326 17.40 0.5385 0.0346 16.71 0.4792 0.0377 17.06 0.4876 0.0350
GridDehazeNet[27] 22.77 0.6145 0.0069 22.62 0.6468 0.0083 23.96 0.7112 0.0061 23.12 0.6575 0.0071
4KDehazing[10] 22.83 0.6177 0.0063 22.47 0.6505 0.0083 22.55 0.6912 0.0099 22.62 0.6531 0.0082
SCAnet[32] 18.37 0.4718 0.0200 18.11 0.0538 0.0210 19.07 0.5966 0.0180 18.52 0.3741 0.0195
Trinity-Net[1] 20.51 0.5728 0.0120 22.46 0.5728 0.0085 24.11 0.7234 0.0058 22.36 0.6230 0.0060
DehazeFormer[7] 22.74 0.6005 0.0071 23.06 0.6137 0.0076 24.69 0.7143 0.0051 23.50 0.6428 0.0066
PCSformer[31] 21.83 0.5427 0.0088 22.09 0.5984 0.0092 23.71 0.6547 0.0055 22.54 0.5996 0.0070
MCAF-Net 23.32 0.6236 0.0059 23.60 0.6583 0.0063 25.40 0.7221 0.0040 24.11 0.6680 0.0054
@@ V-B1 Real-World RSID
Table II presents a detailed performance comparison of the proposed MCAF-Net against several SOTA dehazing methods on the RRSHID dataset, which encompasses real-world RS images under varying haze conditions: thin, moderate, and thick. The results, as shown in the table, indicate that MCAF-Net consistently achieves the highest performance across all haze levels and metrics, marked in red, with the second-best performances highlighted in blue.
RRSHID-thin: MCAF-Net achieves a PSNR of 23.32, an SSIM of 0.6236, and an MSE of 0.0059, surpassing all competing methods. The second-best performer, 4KDehazing [10], records a PSNR of 22.83 and an SSIM of 0.6177, trailing MCAF-Net by 0.49 in PSNR and 0.0059 in SSIM, with a higher MSE of 0.0063.
RRSHID-moderate: MCAF-Net excels with a PSNR of 23.60, an SSIM of 0.6583, and an MSE of 0.0063. In comparison, DehazeFormer [7], the second-best method, achieves a PSNR of 23.06 and an SSIM of 0.6137, falling short by 0.54 in PSNR and 0.0446 in SSIM, with an MSE of 0.0076.
RRSHID-thick: MCAF-Net demonstrates superior performance with a PSNR of 25.40, an SSIM of 0.7221, and an MSE of 0.0040. DehazeFormer obtains a PSNR of 24.69 and an SSIM of 0.7143, lagging by 0.71 in PSNR and 0.0078 in SSIM, with an MSE of 0.0051.
RRSHID-average: Across all haze levels, MCAF-Net maintains its lead with an average PSNR of 24.11, an SSIM of 0.6680, and an MSE of 0.0054. The second-best results vary by metric: DehazeFormer achieves a PSNR of 23.50 (0.61 lower than MCAF-Net), GridDehazeNet [27] records an SSIM of 0.6575 (0.0105 lower), and Trinity-Net [1] yields an MSE of 0.0060 (0.0006 higher).
@@
Compared to established methods such as DCP [8], FFA-Net [12], SCAnet [32] and PCSformer [31], MCAF-Net demonstrates substantial improvements, especially in SSIM and MSE, reflecting its enhanced ability to preserve structural fidelity and minimize reconstruction errors. Even against more recent approaches such as GridDehazeNet, 4KDehazing, Trinity-Net, and DehazeFormer, MCAF-Net stands out as a superior solution, reinforcing its position as a SOTA method for real-world RSID.
To provide a more comprehensive assessment beyond numerical metrics, we included visual comparisons of the dehazing results in Fig. 7. These qualitative examples offered insight into the perceptual improvements achieved by our method. Qualitative results further validate the quantitative findings, showcasing our algorithm’s ability to effectively remove haze while preserving crucial image details and enhancing overall visual clarity, while FFA-Net and Trinity-Net often leave residual haze or introduce artifacts.
@@ V-B2 Synthetic RSID
TABLE III: Performance comparison of different methods on multiple haze datasets. Values in red and blue denote the best and second-best performance, respectively.
@@
PSNR SSIM MSE PSNR SSIM MSE PSNR SSIM MSE PSNR SSIM MSE
DCP[8] 13.87 0.6892 0.0538 15.68 0.6860 0.0491 14.10 0.3898 0.0656 17.87 0.8481 0.0169
FFA-Net[12] 18.31 0.8582 0.0204 23.73 0.9068 0.0096 17.77 0.6261 0.0323 19.45 0.9023 0.0117
GridDehazeNet[27] 23.50 0.9383 0.0080 33.45 0.9766 0.0010 31.54 0.8839 0.0014 20.51 0.9097 0.0089
4KDehazing[10] 23.61 0.9415 0.0053 27.54 0.9425 0.0024 25.21 0.8604 0.0087 20.75 0.7696 0.0085
Trinity-Net[1] 23.60 0.9322 0.0060 23.46 0.8796 0.0100 18.81 0.7781 0.0200 20.43 0.8056 0.0086
PSMB-Net[28] 25.64 0.9447 0.0030 31.32 0.9444 0.0006 30.21 0.8756 0.0010 21.55 0.8489 0.0084
DehazeFormer[7] 25.04 0.9393 0.0040 36.15 0.9794 0.0005 34.54 0.8875 0.0010 22.02 0.9306 0.0063
PCSformer[31] 23.03 0.9159 0.0063 35.17 0.9564 0.0006 33.80 0.8781 0.0011 20.20 0.8125 0.0097
MCAF-Net 25.89 0.9531 0.0032 35.79 0.9814 0.0005 34.59 0.8948 0.0009 22.16 0.9308 0.0062
@@
Our method demonstrated exceptional performance in real-world RSID. To further evaluate its generalization capability, we applied the proposed approach to several publicly available synthetic RSID datasets. The methods include DCP [8], FFA-Net [12], GridDehazeNet [27], 4KDehazing [10], Trinity-Net [1], PSMB-Net [28], DehazeFormer [7] and PCSformer [31] and our proposed MCAF-Net.
@@
Experimental results, presented in Table III, clearly demonstrated that MCAF-Net outperformed nearly all evaluated approaches in terms of quantitative performance on synthetic RS hazy degradation datasets. Notably, our method exhibited particularly high performance in terms of SSIM scores, indicating its superiority in preserving structural information.
@@
V-C Ablation Studies
V-C1 Individual Components
In this study, we employed a baseline architecture with a U-shaped design, utilizing ResNet-Block (RNB) [54] as the basic learning block. Table  IV depicts the ablation on individual components. Initially, we replaced the RNB with our MFIBA, resulting in a performance gain of 0.95 dB in PSNR. Remarkably, this improvement was achieved with only 46.43% #Param and 38.65% FLOPs of the baseline model, demonstrating the efficiency and effectiveness of our MFIBA. Subsequently, incorporating our proposed CSAM module, the model achieved an additional performance gain of 0.46 dB, albeit with a slight increase in computational burden (582.3K #Param and 20.15G FLOPs). This result indicated the capability of our CSAM in improving performance and demonstrated the significance of color correction in RSID. Finally, we integrated our MFAFM, which leaded to a substantial improvement of 0.79 dB in PSNR while maintaining similar computational efficiency as MFIBA (558.1K #Param and 19.82G FLOPs). In conclusion, the ablation studies on individual components demonstrated the significant impact of each proposed component in a positive manner.
TABLE IV: Ablation study of different components.
@@ V-C1 Individual Components
TABLE V: Comparison of different basic learning blocks.
@@
V-C2 Basic Learning Blocks
To assess the effectiveness of our proposed MFIBA, we conducted studies comparing its performance with two commonly used learning blocks, namely RNB [54] and the FasterNetBlock (FNB) [55], in the context of RSID. Table V presented the results of our experiments. Notably, our MFIB×$\times$3 outperformed RNB by a margin of 0.58 dB and FNB by a margin of 0.45 dB in terms of PSNR. Furthermore, we investigated the impact of cascading multiple MFIBs. As shown in Table V, the performance gradually improved as the number of MFIBs increases, with MFIB×$\times$3 achieving the best performance while maintaining relatively low computational cost (558.1K #Param and 19.82G FLOPs). These findings demonstrated the effectiveness of our MFIBA design and its potential for enhancing RSID performance.
V-C3 Capacity of CSAM
To comprehensively evaluate the effectiveness of the CSAM, we conducted detailed ablation studies focusing on its two core components: the self-supervised color correction mechanism and the attention-guided feature refinement, with the results summarized in Table VI.
TABLE VI: Performance Comparison of Different Components

### v2

@@ IV-B Haze Density Estimation
where $I$ is the hazy image, $A$ is the atmospheric light, and $\omega$ is a constant (typically 0.95) to retain subtle haze effects. For each image, we calculated the mean dark channel value as a proxy for haze density. The threshold determination for haze categorization followed a data-driven approach. We performed k-means clustering (k=3) on the distribution of mean dark channel values across the entire dataset, which revealed natural groupings at 110.58 and 159.31. These thresholds were further validated through visual assessment by remote sensing experts and correlation with meteorological visibility data from ground stations (Pearson’s r = 0.87), as depicted in Fig. 6. Images were accordingly categorized into thin (below 110.58), moderate (110.58–159.31), and thick (above 159.31) haze levels.
@@ IV-C Dataset Characteristics
TABLE II: Distribution of Image Pairs in the RRSHID Dataset
@@
RRSHID encompasses 3,053 pairs of hazy and haze-free RS images, constituting the largest real-world dataset of its kind, as comparison in Table I. These pairs are stratified by haze density into three categories, as detailed in Table II.
· Color deviations caused by variations in natural lighting conditions and atmospheric composition.
These properties, as depicted in Fig. 7, render RRSHID a crucial resource for developing robust RSID algorithms, assessing haze impacts on applications such as object detection [57, 58] and change detection [59, 60], and deepening insights into atmospheric effects on satellite imagery.
Figure 7: Sample of the RRSHID that we proposed. Unlike synthetic datasets (e.g., RSID, RS-HAZE), RRSHID demonstrates authentic characteristics in haze density distribution, land cover diversity, and color distortion.
@@ V-A1 Datasets
MCAF-Net is evaluated on the RRSHID dataset and five publicly available benchmarks: RSID [1], RICE1 [10], RICE2 [10], StateHaze1K-thick [54], and HRSI100. These datasets are selected for their diverse characteristics:
@@
HRSI100: Contains 100 real-world RS hazy images cropped to $256\times 256$ resolution from the HRSI [56] dataset without ground truth, specifically designed to evaluate generalization capability in real-world scenarios.
The RRSHID dataset, designed to reflect real-world RS challenges, complements these synthetic and real-world benchmarks to comprehensively assess generalization performance.
@@ V-A3 Evaluation Metrics
To assess the performance on supervised hazy datasets, we employed four metrics: PSNR, Structural Similarity Index Measure (SSIM), Mean Squared Error (MSE), and Learned Perceptual Image Patch Similarity (LPIPS). Higher PSNR and SSIM values signify improved image quality, while lower MSE and LPIPS indicates enhanced dehazing efficacy. To further validate generalization capability in real-world scenarios, we additionally utilized Natural Image Quality Evaluator (NIQE) [61] and Integrated Local NIQE (IL-NIQE) [62] metrics, where lower values indicate better perceptual quality without requiring reference images.
@@ V-B Experimental Results
TABLE III: Performance comparison of different methods on RRSHID dataset across various haze levels. Values in red and blue denote the best and second-best performance, respectively.
@@
PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS
DCP[7] 18.46 0.4564 0.0192 0.4851 17.80 0.4856 0.0238 0.4700 18.39 0.4843 0.0208 0.4996 18.22 0.4754 0.0213 0.4849
FFA-Net [11] 17.08 0.4452 0.0326 0.5761 17.40 0.5385 0.0346 0.5450 16.71 0.4792 0.0377 0.5573 17.06 0.4876 0.0350 0.5595
GridDehazeNet[25] 22.77 0.6145 0.0069 0.4123 22.62 0.6468 0.0083 0.3833 23.96 0.7112 0.0061 0.3947 23.12 0.6575 0.0071 0.3968
4KDehazing[9] 22.83 0.6177 0.0063 0.4352 22.47 0.6505 0.0083 0.4590 22.55 0.6912 0.0099 0.4754 22.62 0.6531 0.0082 0.4565
SCAnet[34] 18.37 0.4718 0.0200 0.4827 18.11 0.0538 0.0210 0.4962 19.07 0.5966 0.0180 0.4734 18.52 0.3741 0.0195 0.4841
Trinity-Net[1] 20.51 0.5728 0.0120 0.4578 22.46 0.5728 0.0085 0.4314 24.11 0.7234 0.0058 0.4206 22.36 0.6230 0.0060 0.4366
DehazeFormer[6] 22.74 0.6005 0.0071 0.4609 23.06 0.6137 0.0076 0.4495 24.69 0.7143 0.0051 0.4377 23.50 0.6428 0.0066 0.4494
PCSformer[31] 21.83 0.5427 0.0088 0.4963 22.09 0.5984 0.0092 0.5021 23.71 0.6547 0.0055 0.4996 22.54 0.5996 0.0070 0.4993
PhDnet[63] 22.64 0.6054 0.0072 0.4658 22.92 0.6448 0.0075 0.4019 24.28 0.6996 0.0053 0.3993 23.28 0.6499 0.0067 0.4223
MCAF-Net 23.32 0.6236 0.0059 0.4023 23.60 0.6583 0.0063 0.3799 25.40 0.7221 0.0040 0.3942 24.11 0.6680 0.0054 0.3921
@@ V-B1 Real-World RSID
Table III presents a detailed performance comparison of the proposed MCAF-Net against several SOTA dehazing methods on the RRSHID dataset, which encompasses real-world RS images under varying haze conditions: thin, moderate, and thick. The results, as shown in the table, indicate that MCAF-Net consistently achieves the highest performance across all haze levels and metrics, marked in red, with the second-best performances highlighted in blue.
RRSHID-thin: MCAF-Net achieves a PSNR of 23.32, an SSIM of 0.6236, an MSE of 0.0059 and a LPIPS of 0.4023, surpassing all competing methods. The second-best performer, 4KDehazing [9], records a PSNR of 22.83 and an SSIM of 0.6177, trailing MCAF-Net by 0.49 in PSNR and 0.0059 in SSIM, with a higher MSE of 0.0063 and LPIPS of 0.4352.
RRSHID-moderate: MCAF-Net excels with a PSNR of 23.60, an SSIM of 0.6583, an MSE of 0.0063 and a LPIPS of 0.3799. In comparison, DehazeFormer [6], the second-best method, achieves a PSNR of 23.06 and an SSIM of 0.6137, falling short by 0.54 in PSNR and 0.0446 in SSIM, with an MSE of 0.0076 and a LPIPS of 0.4495.
RRSHID-thick: MCAF-Net demonstrates superior performance with a PSNR of 25.40, an SSIM of 0.7221, an MSE of 0.0040 and a LPIPS of 0.3942. DehazeFormer obtains a PSNR of 24.69 and an SSIM of 0.7143, lagging by 0.71 in PSNR and 0.0078 in SSIM, with an MSE of 0.0051 and a LPIPS of 0.4377.
RRSHID-average: Across all haze levels, MCAF-Net maintains its lead with an average PSNR of 24.11, an SSIM of 0.6680, an MSE of 0.0054 and a LPIPS of 0.3921. The second-best results vary by metric: DehazeFormer achieves a PSNR of 23.50 (0.61 lower than MCAF-Net), GridDehazeNet [25] records an SSIM of 0.6575 (0.0105 lower) and a LPIPS of 0.3968 (0.0047 higher), and Trinity-Net [1] yields an MSE of 0.0060 (0.0006 higher).
@@
Compared to established methods such as DCP [7], FFA-Net [11], SCAnet [34] and PCSformer [31], MCAF-Net demonstrates substantial improvements, especially in SSIM and MSE, reflecting its enhanced ability to preserve structural fidelity and minimize reconstruction errors. Even against more recent approaches such as GridDehazeNet, 4KDehazing, Trinity-Net, DehazeFormer and PhDnet[63], MCAF-Net stands out as a superior solution, reinforcing its position as a SOTA method for real-world RSID.
To provide a more comprehensive assessment beyond numerical metrics, we included visual comparisons of the dehazing results in Fig. 8. These qualitative examples offered insight into the perceptual improvements achieved by our method. Qualitative results further validate the quantitative findings, showcasing our algorithm’s ability to effectively remove haze while preserving crucial image details and enhancing overall visual clarity, while FFA-Net and 4KDehazing often leave residual haze or introduce artifacts. Despite MCAF-Net’s overall effectiveness across various haze conditions, it still faces challenges in extremely dense haze scenarios. When visibility is severely compromised, the model struggles to recover fine-grained details and occasionally produces artifacts in regions with complex textures. As demonstrated in the last row of Fig. 8, our method shows performance limitations when dealing with intricate structures such as dense urban areas or complex vegetation patterns.
@@ V-B2 Synthetic RSID
TABLE IV: Performance comparison of different methods on multiple haze datasets. Values in red and blue denote the best and second-best performance, respectively.
@@
PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS PSNR SSIM MSE LPIPS
DCP[7] 13.87 0.6892 0.0538 0.2615 15.68 0.6860 0.0491 0.3766 14.10 0.3898 0.0656 0.6234 17.87 0.8481 0.0169 0.1879
FFA-Net[11] 18.31 0.8582 0.0204 0.1449 23.73 0.9068 0.0096 0.1885 17.77 0.6261 0.0323 0.5378 19.45 0.9023 0.0117 0.2357
GridDehazeNet[25] 23.50 0.9383 0.0080 0.0541 33.45 0.9766 0.0010 0.0405 31.54 0.8839 0.0014 0.1942 20.51 0.9097 0.0089 0.1361
4KDehazing[9] 23.61 0.9415 0.0053 0.0693 27.54 0.9425 0.0024 0.0967 25.21 0.8604 0.0087 0.3704 20.75 0.7696 0.0085 0.1674
Trinity-Net[1] 23.60 0.9322 0.0060 0.0692 23.46 0.8796 0.0100 0.0512 18.81 0.7781 0.0200 0.2287 20.43 0.8056 0.0086 0.1764
PSMB-Net[28] 25.64 0.9447 0.0030 0.0537 31.32 0.9444 0.0006 0.0428 30.21 0.8756 0.0010 0.1983 21.55 0.8489 0.0084 0.1642
DehazeFormer[6] 25.04 0.9393 0.0040 0.0557 36.15 0.9794 0.0005 0.0460 34.54 0.8875 0.0010 0.2035 22.02 0.9306 0.0063 0.1707
PCSformer[31] 23.03 0.9159 0.0063 0.0718 35.17 0.9564 0.0006 0.0421 33.80 0.8781 0.0011 0.2418 20.20 0.8125 0.0097 0.1734
PhDnet[63] 25.15 0.9406 0.0040 0.0612 35.50 0.9578 0.0005 0.0464 33.44 0.8877 0.0011 0.2277 21.99 0.8294 0.0064 0.1721
MCAF-Net 25.89 0.9531 0.0032 0.0421 35.79 0.9814 0.0005 0.0383 34.59 0.8948 0.0009 0.1964 22.16 0.9308 0.0062 0.1453
@@
Our method demonstrated exceptional performance in real-world RSID. To further evaluate its generalization capability, we applied the proposed approach to several publicly available synthetic RSID datasets. The methods include DCP [7], FFA-Net [11], GridDehazeNet [25], 4KDehazing [9], Trinity-Net [1], PSMB-Net [28], DehazeFormer [6], PCSformer [31], PhDnet[63] and our proposed MCAF-Net.
@@
Experimental results, presented in Table IV, clearly demonstrated that MCAF-Net outperformed nearly all evaluated approaches in terms of quantitative performance on synthetic RS hazy degradation datasets. Notably, our method exhibited particularly high performance in terms of SSIM scores, indicating its superiority in preserving structural information.
@@
V-B3 The generalization verification in Real-World RSID
Figure 10: Visual samples generated by the evaluated methods on the HRSI100 dataset. Our MCAF-Net successfully eliminates haze and better restores image details, outperforming other evaluated methods.
TABLE V: Generalization comparison of state-of-the-art methods
        Method         NIQE         IL-NIQE
        DCP [7]         6.618         65.79
        FFA-Net [11]         18.609         120.47
        4KDehazing [9]         9.207         91.79
        DehazeFormer [6]         16.282         131.61
        PCSformer [31]         6.105         85.04
        PhDnet [63]         9.684         97.12
        MCAF-Net         5.607         58.32
To further validate the generalization capability of our proposed method and dataset in real-world scenarios, we conducted additional experiments on the HRSI100 dataset. For fair comparison, all methods were evaluated using models trained on the RRSHID-thick dataset without any fine-tuning. Table  V presents the quantitative results, showing that our method achieves the best NIQE and IL-NIQE scores (5.607 and 58.32 respectively), significantly outperforming other comparison methods. This demonstrates superior generalization ability on unseen real-world hazy scenes. Fig. 10 provides visual comparisons that further support our findings - our method not only effectively removes haze but also better preserves image details. These results confirm the effectiveness and generalization capability of our proposed method and dataset in real-world RS applications, which is crucial for practical deployment.
TABLE VI: Ablation study of different components.
@@ V-B3 The generalization verification in Real-World RSID
V-C Ablation Studies
V-C1 Individual Components
In this study, we employed a baseline architecture with a U-shaped design, utilizing ResNet-Block (RNB) [64] as the basic learning block. Table  VI depicts the ablation on individual components. Initially, we replaced the RNB with our MFIBA, resulting in a performance gain of 0.95 dB in PSNR. Remarkably, this improvement was achieved with only 46.43% #Param and 38.65% FLOPs of the baseline model, demonstrating the efficiency and effectiveness of our MFIBA. Subsequently, incorporating our proposed CSAM module, the model achieved an additional performance gain of 0.46 dB, albeit with a slight increase in computational burden (582.3K #Param and 20.15G FLOPs). This result indicated the capability of our CSAM in improving performance and demonstrated the significance of color correction in RSID. Finally, we integrated our MFAFM, which led to a substantial improvement of 0.79 dB in PSNR while maintaining similar computational efficiency as MFIBA (558.1K #Param and 19.82G FLOPs). In conclusion, the ablation studies on individual components demonstrated the significant impact of each proposed component in a positive manner.
TABLE VII: Comparison of different basic learning blocks.
@@ V-C1 Individual Components
TABLE VIII: Performance Comparison of Different Components
@@
V-C2 Basic Learning Blocks
To assess the effectiveness of our proposed MFIBA, we conducted studies comparing its performance with two commonly used learning blocks, namely RNB [64] and the FasterNetBlock (FNB) [65], in the context of RSID. Table VII presented the results of our experiments. Notably, our MFIB×$\times$3 outperformed RNB by a margin of 0.58 dB and FNB by a margin of 0.45 dB in terms of PSNR. Furthermore, we investigated the impact of cascading multiple MFIBs. As shown in Table VII, the performance gradually improved as the number of MFIBs increases, with MFIB×$\times$3 achieving the best performance while maintaining relatively low computational cost (558.1K #Param and 19.82G FLOPs). These findings demonstrated the effectiveness of our MFIBA design and its potential for enhancing RSID performance.